OpenStudy (anonymous):

Solve for the variable 7 = -13 + d

OpenStudy (anonymous):

20

OpenStudy (anonymous):

(x2y)3

OpenStudy (uri):

Lol tell her how to do. :P

OpenStudy (anonymous):

who?

OpenStudy (uri):

7 = -13 + d -13 will shift after = and sign will change so. 7+13 =d 20 =d

OpenStudy (anonymous):

7 = -13 + d +13 +13 7+13=d 20=d